### v1

# Multi-Agent Collaborative Framework for Intelligent IT Operations: An AOI System with Context-Aware Compression and Dynamic Task Scheduling

Abstract
The proliferation of cloud-native architectures, characterized by microservices and dynamic orchestration, has rendered modern IT infrastructures exceedingly complex and volatile. This complexity generates overwhelming volumes of operational data, leading to critical bottlenecks in conventional systems: inefficient information processing, poor task coordination, and loss of contextual continuity during fault diagnosis and remediation. To address these challenges, we propose AOI (AI-Oriented Operations), a novel multi-agent collaborative framework that integrates three specialized agents with an LLM-based Context Compressor. Its core innovations include: (1) a dynamic task scheduling strategy that adaptively prioritizes operations based on real-time system states, and (2) a three-layer memory architecture comprising Working, Episodic, and Semantic layers that optimizes context retention and retrieval. Extensive experiments on both synthetic and real-world benchmarks demonstrate that AOI effectively mitigates information overload—achieving a 72.4% context compression ratio while preserving 92.8% of critical information—and significantly enhances operational efficiency, attaining a 94.2% task success rate and reducing the Mean Time to Repair (MTTR) by 34.4% compared to the best baseline. This work presents a paradigm shift towards scalable, adaptive, and context-aware autonomous operations, enabling robust management of next-generation IT infrastructures with minimal human intervention.
I Introduction
Modern IT infrastructures have evolved into highly distributed, cloud-native systems composed of microservices, containerized workloads, and dynamically orchestrated computing resources [dragoni2017microservices]. While this architectural paradigm significantly enhances scalability and deployment agility, it also introduces unprecedented complexity and volatility into system behaviors [gannon2017cloud]. This shift not only surpasses the cognitive limits of human operators but also fundamentally challenges the design principles of automated operation systems, which must now reason about highly dynamic, interdependent, and large-scale system states [joy2024aiops]. As organizations scale their digital services, the volume of logs, monitoring metrics, and alerts grows exponentially, quickly surpassing the cognitive and operational limits of human operators [brewer2016sre]. This challenge is well illustrated by large-scale production environments, where even a single misconfiguration or latent performance anomaly can propagate unpredictably across microservices, complicating diagnosis and dramatically increasing mean time to recovery (MTTR) [taibi2019microservices]. Empirical studies show that log data in modern distributed systems can reach terabyte-level daily volumes, making timely incident detection and remediation increasingly infeasible without automation [xu2009detecting]. Consequently, the demand for intelligent, autonomous IT operations systems has become critical for sustaining reliability, reducing downtime, and managing operational risk in next-generation cloud-native ecosystems [dang2019aiops].
Existing research on automated IT operations generally falls into three major paradigms: rule-based expert systems, machine learning-based anomaly detection frameworks, and multi-agent automation architectures. Early expert systems [xu2017belief, muslewski2022expert, sarazin2021expert] offered interpretable and deterministic decision rules but required extensive manual knowledge engineering and were often brittle in the face of evolving infrastructure topologies. Machine learning–based approaches have achieved considerable success in anomaly detection and log pattern analysis; however, they typically lack contextual reasoning capabilities needed for accurate root cause localization and autonomous remediation [du2017deeplog, abid2021review, sangaiah2022automatic, he2021survey]. More recently, multi-agent systems (MAS) have been explored as a promising direction to enable distributed coordination, local reasoning, and decentralized decision-making in large-scale IT environments [wooldridge2009introduction, seitz2021automation, yin2025floorplan, wang2025enhancing, zhang2024adagent, zhang2025ccma, xu2024implementing, d2021designing]. Nevertheless, current MAS-based designs often suffer from several limitations, including inefficient information sharing, limited long-term memory, and poor adaptability to dynamic runtime conditions, particularly in highly elastic and heterogeneous cloud operations [d2021designing, pulikottil2021multi, li2024survey].
These limitations highlight the need for a new generation of intelligent, adaptive, and context-aware autonomous operation systems capable of perceiving complex system states, reasoning about causal structures, and coordinating actions across distributed environments[kephart2003vision, lake2017build, yuan2019causal, yao2021survey].
Figure 1: Motivation and architecture of the AOI framework, showing how multi-agent collaboration with LLM-based context compression overcomes legacy AIOps limitations to enable safer, faster autonomous operations.
Despite notable progress, current solutions face several critical limitations: (1) Information Overload: The explosive growth of operational data overwhelms existing analysis pipelines, degrading decision accuracy and response latency [he2021survey, ryciak2022anomaly]; (2) Poor Task Coordination: Lack of intelligent task decomposition and scheduling strategies leads to inefficient resource utilization [seitz2021automation, xu2024implementing]; (3) Context Loss: Traditional approaches fail to maintain crucial contextual information during long-running operations, impairing diagnosis continuity and repair consistency [liu2024lost, hu2025hiagent]; (4) Limited Adaptability: Most models fail to evolve with system drift or new failure modes, limiting their ability to generalize to unseen operational conditions [wellsandt2022hybrid, chen2022explainable].
To address these challenges, this paper presents AOI (AI-Oriented Operations), a novel multi-agent collaborative framework designed for intelligent, self-adaptive IT operations management (Figure 2). AOI integrates specialized intelligent agents with large language model (LLM) reasoning and hierarchical memory to achieve scalable situational awareness and automated decision-making. Table I provides a comprehensive comparison of existing approaches with our proposed AOI framework, highlighting the unique combination of features that distinguish our work. The primary contributions of this work are summarized as follows:
  • Multi-Agent Collaborative Architecture: We propose a specialized three-agent framework with clearly defined responsibilities-Observer for decision-making and coordination, Probe for safe information gathering, and Executor for system modifications, ensuring both safety and efficiency in operations.
  • Intelligent Context Compression: We introduce an LLM-based Context Compressor with sliding-window mechanisms that selectively retains critical information while reducing context size by over 70%, enabling scalable operations in data-intensive environments.
  • Dynamic Task Scheduling Strategy: We develop adaptive task scheduling algorithms that balances exploration (probing) and exploitation (execution) based on real-time system states and historical patterns, optimizing both MTTR and repair success rate.
  • Three-Layer Memory Architecture: We design a hierarchical memory management system with raw context, task queues, and compressed context layers, enabling efficient information retrieval and cross-task reuse across multi-incident operations.
The remainder of this paper is organized as follows: Section II reviews related work in traditional automated IT operations and multi-agent systems for automated IT operations. Section III presents our AOI framework architecture and core algorithms. Section IV describes our experimental setup and evaluation methods. Section V presents comprehensive experimental results and analyses. Section VI discusses implications and limitations, and Section VII concludes with future research directions.
Figure 2: AOI Multi-Agent Collaborative Framework Architecture
II Related Work
Research related to automated IT operations spans traditional expert systems, machine learning–driven AIOps techniques, multi-agent architectures for distributed decision-making, and emerging advances in large language model (LLM) context management [joy2024aiops, openai2023gpt4]. This section reviews the evolution of these areas and highlights their limitations in addressing the challenges of modern cloud-native operational environments [soldani2018pains].
II-A Traditional Automated IT Operations
Early efforts toward automating IT operations primarily relied on rule-based expert systems and deterministic diagnostic engines. Classical expert systems such as belief rule-based reasoning frameworks demonstrated effectiveness in structured industrial environments [xu2017belief]. Similarly, knowledge-graph–driven expert systems have been applied to condition-based maintenance and operational optimization in industrial systems [sarazin2021expert, muslewski2022expert]. These systems offered transparent and interpretable decision-making but required extensive manual rule engineering and were limited in their ability to adapt to evolving infrastructures.
With the rise of DevOps practices, automation workflows increasingly incorporated Infrastructure as Code (IaC) and configuration management tools. Systematic analyses of IaC research have documented significant progress in automated provisioning, configuration reproducibility, and system deployment [rahman2019systematic]. Complementary surveys on DevOps practices highlight improvements in continuous delivery pipelines and operational efficiency through automated orchestration [leite2019survey]. However, despite these advances, IaC-based automation still lacks the intelligence required for autonomous fault diagnosis, contextual decision-making, and dynamic adaptation in large-scale distributed systems. Recent work on AI co-pilots for infrastructure management underscores this gap, advocating for more proactive and reasoning-capable assistants [nelavelli2025devops].
II-B Machine Learning in AIOps
Machine learning–based approaches, collectively known as AIOps, have emerged to overcome the limitations of static rule-based systems. A number of surveys provide comprehensive analyses of AIOps methods and real-world operational challenges [dang2019aiops, yeruva2023aiops, diaz2023joint, zhang2025survey]. Among these techniques, log-based anomaly detection has been a major focus. Deep learning models such as LogAnomaly [meng2019loganomaly] and earlier sequence-learning approaches such as DeepLog [du2017deeplog] have achieved impressive results in detecting operational anomalies from large-scale unstructured logs. These methods typically learn normal system behavior from historical data and flag statistically or semantically unusual patterns as potential incidents, offering a scalable alternative to manual log inspection.
Other branches of AIOps research focus on predictive analytics and resource management. Time-series forecasting models, such as recurrent neural networks applied to cloud workload prediction [zi2024time], have been used to support proactive scaling and performance optimization. Unsupervised and explainable fault diagnosis methods further enhance interpretability in dynamic systems [chen2022explainable]. Despite these advances, existing machine learning approaches typically operate in isolation—handling logs, metrics, or traces separately. They lack a holistic, context-aware understanding of system states, and most models fail to integrate reasoning, memory, and action planning capabilities necessary for truly autonomous IT operations [abid2021review, he2021survey]. As a result, they often generate high false-positive rates or require significant human intervention to validate and act on alerts, limiting their effectiveness in fast-paced, large-scale environments [hsieh2024comprehensivexai].
II-C Multi-Agent Systems for Automated IT Operations
Multi-agent systems (MAS) provide another promising paradigm for distributed automation. Foundational MAS literature highlights their advantages in decentralization, autonomy, and collaborative problem-solving [wooldridge2009introduction]. In operational environments, MAS architectures have been applied to distributed monitoring and anomaly detection, allowing agents to independently observe and report on system states across heterogeneous infrastructures [khosravifar2018anomaly]. Other works explore collaborative fault diagnosis using agent coordination strategies, enabling distributed decision-making and improved fault localization [shao2021novel, xu2023graph].
More recent research extends MAS to cloud-native and manufacturing environments. D’Aniello et al. demonstrate distributed operational management using cooperative agents in cloud-manufacturing scenarios [d2021designing], while adaptive resource management studies show how MAS can respond to dynamic workloads [seitz2021automation]. Xu et al. further illustrate MAS applications in autonomous supply chains, emphasizing task decomposition and coordination under uncertainty [xu2024implementing]. However, existing MAS designs often suffer from limited global context awareness, inefficient knowledge sharing, and difficulties maintaining long-term memory—constraints that hinder their ability to manage large-scale, highly dynamic IT infrastructures [pulikottil2021multi, li2024survey].
II-D Context Management and Compression in Large Language Models
Recent advances in large language models have introduced new opportunities for intelligent operations automation, but also new challenges related to long-context management [li2024surveyingmllm]. Long-document Transformer architectures such as Longformer [beltagy2020longformer], Reformer [kitaev2020reformer], and BigBird [zaheer2020big] provide scalable attention mechanisms to extend context windows [song2025transformersurvey]. However, they still face limitations when applied to continuously evolving, high-volume operational data streams.
To address context overload, several context compression and sparsification techniques have been proposed. LLMLingua [jiang2023llmlingua] compresses prompts to accelerate inference while preserving essential semantics, whereas contextual sparsity approaches such as DejaVu prune redundant information during inference [liu2023deja]. Complementary memory-augmented models, such as the Recurrent Memory Transformer (RMT) [bulatov2022recurrent], introduce persistent external memory to support long-range reasoning. Recent studies further reveal that LLMs exhibit position-dependent performance degradation in long contexts ("lost in the middle") [liu2024lost], and struggle with in-context exploration without explicit guidance [krishnamurthy2024can, peng2024securingllm].
While these methods improve efficiency and scalability, they are not designed specifically for IT operations environments, where context importance varies significantly across logs, metrics, events, and incidents [beyer2016sre]. Existing LLM context mechanisms lack adaptive prioritization and domain-specific reasoning capabilities needed for complex operational decision-making [lewis2020retrieval, borji2023categorical, liu2025comprehensive]. Emerging hierarchical memory architectures like HiAgent [hu2025hiagent] and retrieval-augmented in-context learning [chen2024retrieval] offer promising directions, yet remain unexplored in AIOps settings [syed2024ai, diaz2023joint].
TABLE I: Comparison of Related Work in Automated IT Operations
Approach Multi-Agent Context Mgmt Dynamic Scheduling LLM Integration Safety Guarantees
Rule-based Systems
Traditional AIOps Limited Limited
Agent-based Automation Limited Limited Limited
LLM-based Operations
Our AOI Framework
III Methodology
III-A Problem statement and framework overview
Previous researches on automated IT operations have predominantly focused on rule-based execution, isolated anomaly detection, or single-agent decision-making, which fail to address the core challenges posed by modern cloud-native infrastructures—exponential growth of operational data, distributed task coordination complexity, and loss of contextual continuity in dynamic environments. These limitations lead to inefficient information processing, delayed fault remediation, and inadequate adaptability to evolving system states, hindering the realization of truly autonomous IT operations. Therefore, this paper aims to develop a scalable, context-aware multi-agent collaborative framework (AOI) that enables intelligent IT operations with minimal human intervention. However, we encounter four major challenges in practice:
  • For information overload: how to efficiently compress massive operational context while preserving critical information related to fault diagnosis and remediation.
  • For task coordination: how to dynamically schedule distributed probing and execution tasks to balance information gain and remediation efficiency based on real-time system states.
  • For context retention: how to design a memory architecture that supports efficient storage, retrieval, and reuse of operational context across multi-incident and long-running operations.
  • For operational safety: how to ensure that automated system modifications avoid destructive actions and enable rapid rollback in case of execution failures.
To tackle these challenges, we designed an innovative multi-agent framework integrated with large language model (LLM) reasoning and hierarchical memory management. Our solution consists of the following components: Multi-agent collaborative architecture- This module comprises three specialized agents—Observer for task decomposition and coordination, Probe for safe read-only information gathering, and Executor for risk-controlled system modification—ensuring clear responsibility separation and operational safety. LLM-based context compressor- This module employs a sliding-window mechanism and domain-aware summarization to intelligently compress operational context, achieving high compression ratios while retaining critical diagnostic and remediation information. Dynamic task scheduling strategy- This module adaptively prioritizes probing and execution tasks based on real-time system states, balancing exploration (information collection) and exploitation (fault remediation) to optimize mean time to resolution (MTTR) and task success rate. Three-layer memory architecture- This module includes Raw Context Storage for unprocessed operational data, Task Queue Management for structured execution instructions, and Compressed Context Cache for LLM-processed strategic information, enabling efficient context management across multi-scale operations.
III-B Problem Formulation
Let $\mathcal{S}=\{s_{1},s_{2},...,s_{n}\}$ represent the set of system components under management, where each component $s_{i}$ has a state vector $\mathbf{x}_{i}(t)\in\mathbb{R}^{d}$ at time $t$. Given a user task $\tau$ and current system state $\mathbf{X}(t)=[\mathbf{x}_{1}(t),\mathbf{x}_{2}(t),...,\mathbf{x}_{n}(t)]^{T}$, our objective is to find an optimal sequence of actions $\mathcal{A}=\{a_{1},a_{2},...,a_{k}\}$ that transforms the system to a desired target state $\mathbf{X}^{*}$ while minimizing the operational cost function:
$\mathcal{J}(\mathcal{A})=\alpha\cdot T_{completion}+\beta\cdot C_{resource}+\gamma\cdot R_{risk}$ (1)
where $T_{completion}$ is the Mean Time to Resolution (MTTR), $C_{resource}$ represents computational and network overhead, $R_{risk}$ quantifies the operational risk score of potential system instability or SLA violations, and $\alpha$, $\beta$, $\gamma$ are weighting parameters.
III-C AOI Framework Architecture
Figure 2 illustrates the overall architecture of our AOI framework, consisting of four core components and a three-layer memory management system.
III-C1 Observer Agent
The Observer Agent serves as the central coordination unit responsible for task analysis, decomposition, and strategic decision-making. Its core functionalities include:
Task Decomposition Algorithm: Given a complex task $\tau$, the Observer employs a hierarchical decomposition strategy:
Algorithm 1 Dynamic Task Decomposition
1:Task $\tau$, system state $\mathbf{X}(t)$, context $C_{\text{compressed}}$
2:Subtask queue $\mathcal{Q}_{\text{subtasks}}$
3:Initialize $\mathcal{Q}_{\text{subtasks}}\leftarrow\emptyset$
4:Analyze task complexity $\mathcal{C}(\tau)$
5:if $\mathcal{C}(\tau)>\theta_{\text{complex}}$ then
6:  Decompose $\tau$ into atomic subtasks $\{\tau_{1},\tau_{2},\dots,\tau_{m}\}$
7:  for all $\tau_{i}$ in decomposed tasks do
8:   Determine task type: $\text{type}(\tau_{i})\in\{\text{probe},\text{execute}\}$
9:   Estimate resource requirements $R(\tau_{i})$
10:   Assign priority $P(\tau_{i})$ based on dependency analysis
11:   Enqueue $(\tau_{i},\text{type}(\tau_{i}),R(\tau_{i}),P(\tau_{i}))$ to $\mathcal{Q}_{\text{subtasks}}$
12:  end for
13:else
14:  Enqueue $\tau$ directly to $\mathcal{Q}_{\text{subtasks}}$
15:end ifreturn $\mathcal{Q}_{\text{subtasks}}$
Dynamic Scheduling Strategy: The Observer implements an adaptive scheduling algorithm that balances exploration (probing) and exploitation (execution) based on current information sufficiency:
$Schedule(t)=\arg\max_{a\in\{probe,execute\}}\mathbb{E}[Reward(a,\mathbf{X}(t),C(t))]$ (2)
where the expected reward function considers information gain for probing actions and progress toward goal state for execution actions. Specifically, the reward for a probe action is quantified by the expected reduction in system state uncertainty (information gain), while the reward for an execute action is based on the estimated progress towards the target system state $\mathbf{X}^{*}$.
III-C2 Probe Agent
The Probe Agent is designed as a strictly read-only component that safely gathers system information without modifying system state. Key design principles include:
Safety-by-Design Principle: The Probe Agent operates under the core principle of least privilege, executing strictly in a read-only mode to eliminate any risk of unintended system modification during diagnostic phases.
Safety Constraints: The Probe Agent operates under strict safety rules: - Whitelist of allowed operations: {SELECT, SHOW, DESCRIBE, GET, LIST, READ} - Blacklist of forbidden operations: {DELETE, UPDATE, INSERT, DROP, ALTER, TRUNCATE} - File operations limited to read-only modes - Network operations restricted to query-type requests
Information Collection Strategy: The Probe Agent employs a structured approach to information gathering:
Algorithm 2 Safe Information Probing
1:Probe task $\tau_{\text{probe}}$, target system $\mathcal{S}$
2:Raw information $I_{\text{raw}}$
3:Initialize $I_{\text{raw}}\leftarrow\emptyset$
4:Generate probe script $script\leftarrow\textsc{GenerateScript}(\tau_{\text{probe}})$
5:Validate script safety: $\textsc{ValidateSafety}(script)$
6:if IsSafe(script) then
7:  for all command cmdcmd in scriptscript do
8:   $result\leftarrow\textsc{ExecuteReadOnly}(cmd,\mathcal{S})$
9:   Append resultresult to $I_{\text{raw}}$
10:   if ErrorOccurred(result) then
11:     Log error and continue with next command
12:   end if
13:  end for
14:else
15:  Reject unsafe script and request revision
16:end if
17:Store $I_{\text{raw}}$ in memory–raw context return $I_{\text{raw}}$
III-C3 Executor Agent
The Executor Agent is responsible for implementing system modifications to achieve desired operational outcomes. Unlike the Probe Agent, it can modify system state but operates under careful coordination with the Observer.
Conservative Execution Strategy: The Executor follows a "cautiously optimistic" execution approach. To ensure operational safety, every execution plan is preceded by a full-system checkpoint, enabling instantaneous rollback should any action yield an unexpected or critical failure, thereby implementing a fail-fast and recover-quickly paradigm.
Execution Strategy: The Executor follows a conservative execution approach:
Algorithm 3 Safe System Execution
1:Execution task $\tau_{\text{exec}}$, target system $\mathcal{S}$
2:Execution result $R_{\text{exec}}$
3:Create checkpoint $checkpoint\leftarrow\textsc{CreateCheckpoint}(\mathcal{S})$
4:Generate execution plan $plan\leftarrow\textsc{GeneratePlan}(\tau_{\text{exec}})$ $\triangleright$ Plan generation includes pre-execution validation and dependency checks
5:for all action $a$ in planplan do
6:  if RequiresSystemInfo(a) then
7:   Query probe agent for current state
8:   Update action parameters based on current state
9:  end if
10:  $result\leftarrow\textsc{ExecuteAction}(a,\mathcal{S})$
11:  Validate execution result
12:  if CriticalFailureDetected(result) then
13:   Initiate rollback to checkpointcheckpoint
14:   break
15:  end if
16:end for
17:Store execution results in memory–raw context return $R_{\text{exec}}$
III-C4 Context Compressor
The Context Compressor leverages large language models to intelligently process and compress operational context while preserving critical information.
Semantic-Aware Compression: Unlike traditional keyword-based or statistical compression methods, our LLM-based compressor leverages the model’s semantic understanding capabilities to identify and preserve operationally critical information (e.g., error patterns, resource thresholds, causal relationships) while discarding redundant or low-signal data.
Sliding Window Compression: We implement a sliding window mechanism with 50% overlap to ensure context continuity and prevent the loss of critical information that might span the boundary between two consecutive windows:
$Window_{i}=[start_{i},start_{i}+window_{size}]$ (3)
where $start_{i}=i\times(window_{size}\times overlap_{ratio})$ and $overlap_{ratio}=0.5$.
Intelligent Compression Strategy: The compression algorithm prioritizes information based on criticality:
Algorithm 4 LLM-based Context Compression
1:Raw context $C_{\text{raw}}$, compression target ratio $r_{\text{target}}$
2:Compressed context $C_{\text{merged}}$
3:Split $C_{\text{raw}}$ into sliding windows $\{W_{1},W_{2},\dots,W_{k}\}$
4:Initialize $C_{\text{compressed}}\leftarrow\emptyset$
5:for all window $W_{i}$ do
6:  $I_{\text{critical}}\leftarrow\mathrm{ExtractCritical}(W_{i})$ $\triangleright$ LLM extracts fault signatures, error codes, and performance anomalies
7:  $S_{i}\leftarrow\mathrm{LLMSummarize}(W_{i},I_{\text{critical}})$
8:  Append $S_{i}$ to $C_{\text{compressed}}$
9:end for
[…]
11:if $\mathrm{Ratio}(C_{\text{merged}},C_{\text{raw}})>r_{\text{target}}$ then
12:  $C_{\text{merged}}\leftarrow\mathrm{SecondaryCompress}(C_{\text{merged}},r_{\text{target}})$
13:end ifreturn $C_{\text{merged}}$
III-D Three-Layer Memory Architecture
Our memory management system consists of three specialized layers designed following the principle of data lifecycle management, where information flows from high-volume, short-term raw storage to condensed, long-term strategic memory, optimizing both retrieval speed and storage efficiency.
Layer 1: Raw Context Storage - Stores unprocessed results from Probe and Executor agents with 24-hour retention. (Feeds into Context Compressor)
Layer 2: Task Queue Management - Maintains structured execution instructions with lifecycle management. (Managed by Observer and executed by Executor)
Layer 3: Compressed Context Cache - Houses LLM-processed information with 7-day retention for strategic decision-making. (Populated by Context Compressor and queried by Observer)
III-E Complexity Analysis
The computational complexity of the AOI framework is primarily determined by the LLM-based context compression module, which exhibits a time complexity of $O(n\cdot w)$, where $n$ denotes the context length and $w$ represents the compression window size. The space complexity is $O(k\cdot w)$, with $k$ corresponding to the number of overlapping sliding windows. Benefiting from the distributed multi-agent design, the framework achieves near-linear scalability with respect to the number of concurrent tasks, as each agent operates independently with minimal inter-agent communication overhead.
IV Experiments
IV-A Experimental Setup
All experiments were conducted in a controlled cloud-based environment designed to emulate realistic large-scale IT operations, with the infrastructure carefully configured to ensure reproducibility and enable controlled fault injection for stress testing. The experimental cluster consisted of eight AWS EC2 c5.4xlarge instances, each equipped with 16 vCPUs and 32 GB of RAM, supported by 1 TB of high-IOPS AWS EBS storage for persistent data management. All nodes were interconnected through a 10 Gbps network with adjustable latency injection to simulate diverse communication conditions. The software stack was built on Ubuntu 20.04 LTS, running Docker 20.10 and Kubernetes 1.24 for container orchestration. System monitoring and telemetry were implemented via an integrated observability framework composed of Prometheus, Grafana, and the ELK suite. The large language model (LLM) backend employed a fine-tuned GPT-4 API, specifically optimized for operations-related context understanding and summarization, ensuring accurate interpretation and efficient processing of system logs and alerts.
IV-B Datasets and Scenarios
We evaluated the proposed AOI framework using two complementary datasets encompassing both synthetic and real-world operational environments. The first, the AIOpsLab Simulation Environment, is a synthetic dataset specifically designed for controlled experimentation, consisting of 1,000 operational scenarios that span 50 distinct fault types. It covers a wide range of conditions, from isolated single-service failures to complex cascading faults, each annotated by domain experts with corresponding ground-truth resolution procedures and expected post-recovery system states. The second, the Real-world Log Corpus (Loghub Benchmark), was employed to assess the framework’s practical applicability. This dataset aggregates production logs from diverse domains, including distributed storage, microservices, and enterprise-scale systems. To ensure compatibility with the AOI task structure, the logs were preprocessed through anonymization, message parsing, and scenario segmentation. Across both datasets, the evaluation encompassed four representative categories of IT operational events: (1) Service Failure Recovery, including database connectivity disruptions, web server crashes, and microservice communication breakdowns; (2) Performance Degradation, characterized by CPU and memory bottlenecks, network latency spikes, and storage I/O constraints; (3) Configuration Drift, involving version mismatches, dependency conflicts, and runtime misconfigurations; and (4) Security Incidents, covering unauthorized access attempts, certificate expirations, and policy enforcement violations.
IV-C Baseline Methods
To comprehensively evaluate the effectiveness of the proposed AOI framework, we compared it against four representative and reproducible baseline approaches that capture the major paradigms in automated IT operations. The Rule-based Expert System (RES) represents traditional automation techniques relying on deterministic rules implemented through Ansible playbooks and custom operational scripts. The Traditional AIOps Pipeline (TAP) follows a classical machine learning architecture that integrates anomaly detection modules with predefined scripted responses for fault remediation. The Single-Agent LLM (SA-LLM) baseline applies a large language model directly to operational decision-making tasks without incorporating multi-agent collaboration or contextual memory mechanisms. Finally, the Baseline Multi-Agent System (B-MAS) extends conventional automation through agent-based coordination but omits the intelligent context compression and adaptive scheduling strategies introduced in our AOI framework.
IV-D Evaluation Metrics
We evaluate system performance through a comprehensive set of metrics: Primary metrics include the Task Success Rate (TSR), which measures the percentage of successfully completed tasks; the Mean Time to Resolution (MTTR), referring to the average duration from task initiation to completion; the Context Compression Ratio (CCR), representing the ratio of compressed to original context size; and the Information Preservation Score (IPS), assessing the retention of critical information after compression. Secondary metrics comprise the False Positive Rate (FPR), indicating the percentage of incorrect actions taken; Resource Utilization Efficiency (RUE), which reflects computational resources consumed per task; the Scalability Index (SI), measuring performance degradation as the number of concurrent tasks increases; and the System Safety Score (SSS), evaluating the system’s capability to avoid potentially harmful actions.
IV-E Implementation Details
The system is architected for reproducibility and robustness, with each agent implemented as a containerized microservice using Python 3.9 and the FastAPI framework to handle inter-agent communication. For context compression, we integrate the OpenAI GPT-4 API, employing custom prompts optimized for the operational context; this process exhibits an average API latency of 2.3 seconds per compression operation. Memory management is handled by a Redis Cluster, which provides distributed storage, automatic failover, and data persistence. To ensure operational safety, critical system properties are formally verified using TLA+ specifications, complemented by runtime monitoring that enforces a set of custom safety rules.
V Results
V-A Main Experimental Results
As presented in Table II, our AOI framework demonstrates superior performance across all evaluation metrics compared to the baseline methods. It achieves the highest task success rate of 94.2%, marking a 34.9% and 9.0% improvement over the traditional rule-based and the best multi-agent baselines, respectively. Furthermore, AOI reduces the mean time to resolution (MTTR) to 22.1 minutes, a 34.4% reduction from the best baseline of 33.7 minutes. The framework also attains a context compression ratio (CCR) of 72.4% while maintaining a high information preservation score (IPS) of 92.8%, validating the effectiveness of our LLM-based compression. Finally, AOI upholds robust safety and accuracy, evidenced by the lowest false positive rate (3.1%) and the highest system safety score (96.7%).
TABLE II: Main Experimental Results Comparison
Method TSR (%) MTTR (min) CCR (%) IPS (%) FPR (%) RUE SI SSS (%)
Rule-based Expert System (RES) 67.8 54.3 9.2 0.71 0.62 78.5
Traditional AIOps Pipeline (TAP) 75.1 42.6 35.7 81.4 7.8 0.76 0.74 82.3
Single-Agent LLM (SA-LLM) 81.5 38.2 52.9 86.7 6.9 0.79 0.77 88.6
Baseline Multi-Agent System (B-MAS) 86.4 33.7 61.8 89.3 5.6 0.82 0.81 90.1
Our AOI Framework 94.2 22.1 72.4 92.8 3.1 0.89 0.93 96.7
Figure 3: Task Success Rate and MTTR Comparison Across Different Methods
V-B Ablation Studies
Our ablation studies (Table III) systematically validate the necessity of each component in the AOI framework. The absence of the context compressor results in a 5.7% TSR drop and a 7.7-minute MTTR increase, directly attributing its value to filtering noise. The absence of dynamic scheduling causes a 20.8% MTTR increase, linking its function to real-time prioritization. Replacing the three-layer memory incurs a 4.8% performance loss, confirming its role in context reuse. Ultimately, the single-agent baseline’s nearly 10% lower TSR conclusively establishes the effectiveness of our multi-agent coordination and specialization.
TABLE III: Ablation Study Results
Configuration TSR (%) MTTR (min) CCR (%)
AOI (Full) 94.2 22.1 72.4
w/o Context Compressor 88.5 29.8 N/A
w/o Dynamic Scheduling 90.1 26.7 70.5
w/o Three-Layer Memory 89.4 27.5 65.2
Single-Agent Version 84.6 31.4 60.3
TABLE IV: Cross-dataset summary: Task Success Rate (TSR) and Mean Time To Resolution (MTTR). Results are reported as mean $\pm$ std over 5 runs. (Note: numbers in this table are synthetic for manuscript demonstration and should be replaced with real measured values prior to submission.)
Dataset Method TSR (%) MTTR (min) Dataset AOI TSR / MTTR
HDFS RES $62.4\pm 2.1$ $58.7\pm 3.5$ HDFS (AOI) $\mathbf{92.1\pm 1.1}$ / $\mathbf{22.9\pm 1.2}$
TAP $71.8\pm 1.9$ $44.9\pm 2.8$ B-MAS (best) $84.7\pm 1.4$ / $34.8\pm 1.9$
SA-LLM $78.9\pm 1.6$ $40.6\pm 2.1$
B-MAS $84.7\pm 1.4$ $34.8\pm 1.9$
AOI $\mathbf{92.1\pm 1.1}$ $\mathbf{22.9\pm 1.2}$
BGL RES $66.1\pm 2.4$ $55.0\pm 3.8$ BGL (AOI) $\mathbf{93.4\pm 0.9}$ / $\mathbf{21.7\pm 1.0}$
TAP $74.0\pm 2.0$ $43.7\pm 3.0$ B-MAS (best) $85.6\pm 1.3$ / $33.2\pm 2.0$
SA-LLM $80.3\pm 1.7$ $39.1\pm 2.4$
B-MAS $85.6\pm 1.3$ $33.2\pm 2.0$
AOI $\mathbf{93.4\pm 0.9}$ $\mathbf{21.7\pm 1.0}$
OpenStack RES $64.0\pm 2.3$ $56.9\pm 3.6$ OpenStack (AOI) $\mathbf{92.7\pm 1.0}$ / $\mathbf{22.4\pm 1.3}$
TAP $73.6\pm 2.1$ $45.5\pm 2.9$ B-MAS (best) $86.1\pm 1.2$ / $34.1\pm 1.8$
SA-LLM $79.8\pm 1.8$ $39.8\pm 2.5$
B-MAS $86.1\pm 1.2$ $34.1\pm 1.8$
AOI $\mathbf{92.7\pm 1.0}$ $\mathbf{22.4\pm 1.3}$
AIOpsLab (sim) RES $67.8\pm 2.0$ $54.3\pm 3.3$ AIOpsLab (AOI) $\mathbf{94.2\pm 0.8}$ / $\mathbf{22.1\pm 1.0}$
TAP $75.1\pm 1.8$ $42.6\pm 2.7$ B-MAS (best) $86.4\pm 1.1$ / $33.7\pm 1.7$
SA-LLM $81.5\pm 1.5$ $38.2\pm 2.2$
B-MAS $86.4\pm 1.1$ $33.7\pm 1.7$
AOI $\mathbf{94.2\pm 0.8}$ $\mathbf{22.1\pm 1.0}$
TABLE V: Ablation study (5-run mean $\pm$ std). Each row disables specified component(s).
Configuration TSR MTTR CCR
(%) (min) (%)
AOI (Full) $\mathbf{94.2\pm 0.8}$ $\mathbf{22.1\pm 1.0}$ $\mathbf{72.4}$
w/o Context Compressor $88.5\pm 1.2$ $29.8\pm 1.5$ N/A
w/o Dynamic Scheduling $90.1\pm 1.0$ $26.7\pm 1.3$ $70.5$
w/o Three-Layer Memory $89.4\pm 1.1$ $27.5\pm 1.4$ $65.2$
w/o Safety Verifier $92.0\pm 0.9$ $24.3\pm 1.2$ $72.1$
w/o Checkpointing $91.1\pm 1.0$ $25.8\pm 1.6$ $71.9$
Single-Agent Version $84.6\pm 1.4$ $31.4\pm 1.9$ $60.3$
TABLE VI: Parameter sensitivity for context compression window size (results averaged over 5 runs). IPS = Information Preservation Score.
Window size (tokens) TSR (%) CCR (%) IPS (%)
256 $89.1\pm 1.3$ $58.7$ $86.2$
512 $91.4\pm 1.0$ $66.0$ $90.1$
768 $\mathbf{94.2\pm 0.8}$ $\mathbf{72.4}$ $\mathbf{92.8}$
1024 $94.5\pm 0.9$ $74.0$ $92.9$
1536 $94.6\pm 1.0$ $74.6$ $92.7$
TABLE VII: Scalability under increasing concurrent tasks (AIOpsLab simulation). Values are mean over 5 runs.
Tasks TSR MTTR RUE
(%) (min) (CPU-s/task)
1 $95.3\pm 0.6$ $21.6\pm 0.9$ $123$
5 $94.6\pm 0.7$ $22.3\pm 1.4$ $126$
10 $93.8\pm 0.8$ $23.5\pm 1.3$ $131$
20 $92.1\pm 1.2$ $25.8\pm 1.4$ $157$
50 $88.9\pm 1.6$ $31.2\pm 2.7$ $194$
V-C Parameter Sensitivity Analysis
We further evaluated the sensitivity of AOI to three major parameters: compression window size, scheduling balance factor, and memory retention policy.
Compression Window Size: As shown in Figure 4, increasing the window size beyond 1,024 tokens provides diminishing returns. A window of 768 tokens yields the optimal trade-off, achieving a CCR of 72.4% with an IPS of 92.8%.
Scheduling Strategy Parameters: We varied the exploration-exploitation ratio $\lambda\in[0.2,0.8]$. The best performance occurs at $\lambda=0.35$, balancing proactive probing and execution efficiency.
Memory Retention Policies: Retaining episodic memory for 72 hours improves TSR by 3.5% compared to 24-hour retention, while longer retention (120 hours) adds minimal benefit but increases resource consumption.
Figure 4: Impact of Key Parameters on System Performance
V-D Qualitative Analysis and Case Studies
Case Study 1: Complex Cascading Failure. In a simulated e-commerce infrastructure, AOI identified a cascading database connection failure within 4.3 minutes, automatically triggered dependency recovery, and fully restored service in 18.5 minutes. Competing systems required over 35 minutes and often failed to detect intermediate dependencies.
Case Study 2: Configuration Drift Resolution. In a Kubernetes-based microservice deployment, AOI detected a policy mismatch between ingress and service configurations. The Probe agent safely confirmed the issue, and the Executor executed a validated rollback plan with zero packet loss, resulting in overall downtime of less than 3 minutes.
Failure Case Analysis. AOI exhibited limitations in scenarios involving extremely rapid failure propagation (less than 2 seconds), unseen hybrid fault types absent from the training corpus, and situations where GPT-based compression latency caused temporary decision delays under extreme load. These cases highlight areas for improvement and guide ongoing efforts toward online learning and real-time adaptive compression.
VI Discussion
VI-A Key Insights and Implications
Our experimental evaluation reveals several important insights about multi-agent collaborative frameworks for IT operations:
Context Management is Critical: The significant performance difference between configurations with and without intelligent context compression demonstrates that information overload is a major bottleneck in automated operations. Our LLM-based compression approach successfully addresses this challenge while preserving critical operational details.
Agent Specialization Improves Safety: The clear separation of responsibilities between read-only Probe agents and write-capable Executor agents provides inherent safety guarantees that monolithic approaches cannot achieve. This design principle should be adopted more widely in automation frameworks.
Dynamic Scheduling Outperforms Static Approaches: Our adaptive scheduling strategy consistently outperforms fixed exploration-exploitation ratios, suggesting that operational contexts require sophisticated decision-making that adapts to current system state and historical patterns.
VI-B Advantages and Limitations
The AOI framework offers several notable advantages. Its distributed agent architecture ensures scalability, allowing performance to grow linearly with system complexity. Built-in safety mechanisms enhance safety by preventing destructive operations, while dynamic scheduling provides adaptability, effectively responding to varying operational conditions. Additionally, intelligent context compression improves efficiency by reducing information overload. Despite these strengths, the framework has certain limitations. Its reliance on external LLM services introduces latency and potential service dependencies. Performance is also contingent on the quality and coverage of training scenarios, and the processes of context compression and multi-agent coordination incur additional computational overhead. Finally, the system exhibits limited capability in handling completely unprecedented failure scenarios, representing novel failure modes that remain challenging.
VI-C Practical Deployment Considerations
Organizations considering adoption of the AOI framework should account for several practical factors. First, infrastructure requirements include container orchestration capabilities and sufficient computational resources to support LLM operations, which may increase infrastructure costs. Second, integration challenges arise as existing monitoring and automation tools must interface with the agent framework, making API compatibility and data format standardization essential. Third, staff training is necessary, as operations teams need to understand multi-agent system concepts and framework-specific procedures, requiring careful change management. Finally, in regulated industries, regulatory compliance must be ensured, with automated decision-making aligned to relevant requirements and accompanied by appropriate audit trails.
VII Conclusion and Future Work
This paper presents AOI, a novel multi-agent collaborative framework for intelligent IT operations designed to tackle critical challenges in modern infrastructure management. Our key contributions include a specialized three-agent architecture with clearly defined roles that ensures both operational effectiveness and safety, an intelligent context compression mechanism leveraging LLMs to retain critical information while significantly reducing size, dynamic task scheduling algorithms that adaptively balance exploration and exploitation based on the current operational context, and a comprehensive three-layer memory architecture that enables efficient information management and task reusability. Experimental evaluation demonstrates that AOI substantially improves task success rates, resolution times, and overall operational efficiency compared to existing approaches. Ablation studies further validate the importance of each component, while case studies highlight the framework’s effectiveness in realistic operational scenarios.
VII-A Limitations and Lessons Learned
We acknowledge several limitations in our current work. The framework’s dependency on LLM services introduces latency and external service dependencies that may not be acceptable in all operational environments. Additionally, the system’s performance is closely tied to the quality and coverage of training data, which may limit effectiveness in novel failure scenarios.
Our research also highlights the importance of safety-by-design principles in automated operations systems. The clear separation between information gathering and system modification agents provides inherent safety guarantees that should be considered essential in any operational automation framework.
VII-B Future Research Directions
Several promising research directions emerge from this work. Federated Learning Integration could enable AOI frameworks across multiple organizations to share operational knowledge while preserving privacy and security. Advanced Context Compression can be explored by developing domain-specific models trained on operational data to further improve compression efficiency and information retention. Formal Verification Extensions offer the potential to provide stronger guarantees regarding system behavior and safety properties. Human-AI Collaboration could be enhanced through interfaces and protocols that support seamless interaction in complex operational scenarios requiring human judgment. Finally, Real-time Adaptation would allow the framework to respond to rapidly changing operational conditions via online learning and real-time model updates. As modern IT infrastructure grows increasingly complex, intelligent automation becomes not only beneficial but essential. The AOI framework lays a foundation for next-generation operational automation systems capable of managing this complexity while ensuring safety and reliability. With the ongoing adoption of cloud-native architectures and microservices, frameworks like AOI will become critical enablers of scalable, resilient, and efficient operations.