### v2

# AOI: Context-Aware Multi-Agent Operations via Dynamic Scheduling and Hierarchical Memory Compression

Abstract
The proliferation of cloud-native architectures, characterized by microservices and dynamic orchestration, has rendered modern IT infrastructures exceedingly complex and volatile. This complexity generates overwhelming volumes of operational data, leading to critical bottlenecks in conventional systems: inefficient information processing, poor task coordination, and loss of contextual continuity during fault diagnosis and remediation. To address these challenges, we propose AOI (AI-Oriented Operations), a novel multi-agent collaborative framework that integrates three specialized agents with an LLM-based Context Compressor. Its core innovations include: (1) a dynamic task scheduling strategy that adaptively prioritizes operations based on real-time system states, (2) a three-layer memory architecture comprising Working, Episodic, and Semantic layers that optimizes context retention and retrieval. Extensive experiments on synthetic and real-world benchmarks show that AOI achieves 72.4% context compression while preserving 92.8% critical information, improves task success to 94.2%, and reduces MTTR by 34.4% over the best baseline. This work presents a paradigm shift towards scalable, adaptive, and context-aware autonomous operations, enabling robust management of next-generation IT infrastructures with minimal human intervention.
AOI: Context-Aware Multi-Agent Operations via Dynamic Scheduling and Hierarchical Memory Compression
Zishan Bai1, Jing Luo2, Ziyi Ni3, Enze Ge4, Jiacheng Shi5, Yichao Zhang6, Jiayi Gu7, Zhimo Han8, Riyang Bao9, Junfeng Hao10, 1Columbia University  2Chongqing University 3Institute of Automation, Chinese Academy of Sciences 4AI Agent Lab, Vokram Group  5College of William and Mary 6University of Texas  7Central University of Finance and Economics 8Zhengzhou University of Light Industry  9Emory University 10Department of Nephrology, Affiliated Hospital of Guangdong Medical University Correspondence: ygzhjf85@gmail.com
1 Introduction
Modern IT infrastructures have evolved into highly distributed, cloud-native systems composed of microservices, containerized workloads, and dynamically orchestrated computing resources Dragoni et al. (2017); Yu (2025); Bao et al. (2026). While this architectural paradigm significantly enhances scalability and deployment agility, it also introduces unprecedented complexity and volatility into system behaviors Gannon et al. (2017). This shift not only surpasses the cognitive limits of human operators but also fundamentally challenges the design principles of automated operation systems, which must now reason about highly dynamic, interdependent, and large-scale system states Joy et al. (2024). As organizations scale their digital services, the volume of logs, monitoring metrics, and alerts grows exponentially, quickly surpassing the cognitive and operational limits of human operators Beyer et al. (2016). This challenge is well illustrated by large-scale production environments, where even a single misconfiguration or latent performance anomaly can propagate unpredictably across microservices, complicating diagnosis and dramatically increasing mean time to recovery (MTTR) Taibi et al. (2019). Empirical studies show that log data in modern distributed systems can reach terabyte-level daily volumes, making timely incident detection and remediation infeasible without automation Xu et al. (2009). Consequently, the demand for intelligent, autonomous IT operations systems has become critical for sustaining reliability, reducing downtime, and managing operational risk in next-generation cloud-native ecosystems Dang et al. (2019).
Figure 1: Motivation and architecture of the AOI framework, showing how multi-agent collaboration with LLM-based context compression overcomes legacy AIOps limitations to enable safer, faster autonomous operations.
Existing research on automated IT operations generally falls into three major paradigms: rule-based expert systems, machine learning-based anomaly detection frameworks, and multi-agent automation architectures. Early expert systems Xu et al. (2020); Muślewski et al. (2022); Sarazin et al. (2021); Yu et al. (2025b, c) offered interpretable and deterministic decision rules but required extensive manual knowledge engineering and were often brittle in the face of evolving infrastructure topologies. Machine learning–based approaches have achieved considerable success in anomaly detection and log pattern analysis; however, they typically lack contextual reasoning capabilities needed for accurate root cause localization and autonomous remediation Du et al. (2017); Abid et al. (2021); Sangaiah et al. (2022); He et al. (2021). More recently, multi-agent systems (MAS) have been explored as a promising direction to enable distributed coordination, local reasoning, and decentralized decision-making in large-scale IT environments Wooldridge (2009); Seitz et al. (2021); Yin et al. (2025); Wang et al. (2025); Zhang et al. (2024, 2025b); Xu et al. (2024); D’Aniello et al. (2021). Nevertheless, current MAS-based designs often suffer from several limitations, including inefficient information sharing, limited long-term memory, and poor adaptability to dynamic runtime conditions, particularly in highly elastic and heterogeneous cloud operations D’Aniello et al. (2021); Pulikottil et al. (2021); Li et al. (2024b). These limitations highlight the need for a new generation of intelligent, adaptive, and context-aware autonomous operation systems capable of perceiving complex system states, reasoning about causal structures, and coordinating actions across distributed environments Kephart and Chess (2003); Lake et al. (2017); Pearl et al. (2016); Yao et al. (2021); Ni et al. (2025a).
Despite progress, existing solutions face key limitations: (1) Information Overload: Rapid growth of operational data overwhelms analysis pipelines, degrading decision accuracy and response latency He et al. (2021); Ryciak et al. (2022); (2) Poor Task Coordination: Lack of intelligent task decomposition and scheduling leads to inefficient resource utilization Seitz et al. (2021); Xu et al. (2024); (3) Context Loss: Traditional approaches fail to preserve crucial context in long-running operations, impairing diagnosis continuity and repair consistency Liu et al. (2024); Hu et al. (2025); (4) Limited Adaptability: Most models cannot evolve with system drift or new failure modes, limiting generalization to unseen conditions Wellsandt et al. (2022); Chen et al. (2022); Ni et al. (2025b).
To address these challenges, we present AOI (AI-Oriented Operations), a novel multi-agent framework for intelligent self-adaptive IT operations management (Figure 1). AOI integrates specialized agents with LLM reasoning and hierarchical memory to enable scalable situational awareness and automated decision-making. Table 1 compares AOI with existing approaches and highlights its distinctive features. Our main contributions are:
Figure 2: AOI Multi-Agent Collaborative Framework Architecture. The framework consists of three specialized agents (Observer, Probe, Executor) coordinated through a central Context Compressor, with a three-layer memory architecture for efficient information management.
  • Multi-Agent Collaborative Architecture: We propose a three-agent framework—Observer (coordination), Probe (safe information gathering), and Executor (system modification)—to ensure safe and efficient operations.
  • Intelligent Context Compression: We introduce an LLM-based context compressor with sliding windows that preserves critical information while reducing context size by over 70
  • Dynamic Task Scheduling: We develop adaptive scheduling that balances probing and execution based on real-time states and historical patterns to optimize MTTR and repair success.
  • Three-Layer Memory Architecture: We design a hierarchical memory with raw context, task queues, and compressed context layers for efficient retrieval and cross-task reuse.
Table 1: Comprehensive Comparison of Related Work in Automated IT Operations
Approach Multi-Agent Context Dynamic LLM Safety Memory Theoretical
Coordination Compression Scheduling Integration Guarantees Hierarchy Foundation
Expert Systems Xu et al. (2020); Sarazin et al. (2021)
DeepLog Du et al. (2017) Limited
LogAnomaly Meng et al. (2019) Limited
Traditional AIOps Dang et al. (2019) Limited Limited
MAS Automation Seitz et al. (2021) Limited Limited
Cloud-MAS D’Aniello et al. (2021) Limited Limited Limited
LLMLingua Jiang et al. (2023) Limited
HiAgent Hu et al. (2025) Limited
AOI (Ours)
2 Related Work
Research related to automated IT operations spans traditional expert systems, machine learning–driven AIOps techniques, multi-agent architectures for distributed decision-making, and emerging advances in large language model (LLM) context management Joy et al. (2024); OpenAI (2023). This section reviews the evolution of these areas and highlights their limitations in addressing the challenges of modern cloud-native operational environments Soldani et al. (2018).
2.1 Traditional Automated IT Operations
Early efforts toward automating IT operations primarily relied on rule-based expert systems and deterministic diagnostic engines. Classical expert systems such as belief rule-based reasoning frameworks demonstrated effectiveness in structured industrial environments Xu et al. (2020). Similarly, knowledge-graph–driven expert systems have been applied to condition-based maintenance and operational optimization in industrial systems Sarazin et al. (2021); Muślewski et al. (2022). These systems offered transparent and interpretable decision-making but required extensive manual rule engineering and were limited in their ability to adapt to evolving infrastructures.
Beyond rule-based expert systems, a line of research has investigated formal reliability and diagnosability properties of complex networked systems, providing theoretical foundations for fault detection and isolation. Prior studies have analyzed conditional diagnosability, connectivity, and fault tolerance of various interconnection networks under rigorous diagnostic models, demonstrating how structural properties directly impact system robustness and recovery guarantees. Such works offer valuable insights into how system topology and fault models influence safe operation and diagnosis, which motivates the safety-aware and reliability-driven design principles adopted in autonomous IT operations frameworks Wang et al. (2017b); Wang and Wang (2021); Wang et al. (2024, 2017a).
With the rise of DevOps practices, automation workflows increasingly incorporated Infrastructure as Code (IaC) and configuration management tools. Systematic analyses of IaC research have documented significant progress in automated provisioning, configuration reproducibility, and system deployment Rahman et al. (2019). Complementary surveys on DevOps practices highlight improvements in continuous delivery pipelines and operational efficiency through automated orchestration Leite et al. (2019). However, despite these advances, IaC-based automation still lacks the intelligence required for autonomous fault diagnosis, contextual decision-making, and dynamic adaptation in large-scale distributed systems. Recent work on AI co-pilots for infrastructure management underscores this gap, advocating for more proactive and reasoning-capable assistants Yang et al. (2025).
2.2 Machine Learning in AIOps
Machine learning–based approaches, collectively known as AIOps, have emerged to overcome the limitations of static rule-based systems. A number of surveys provide comprehensive analyses of AIOps methods and real-world operational challenges Dang et al. (2019); Yeruva and Ramu (2023); Diaz-De-Arcaya et al. (2023); Zhang et al. (2025a). Among these techniques, log-based anomaly detection has been a major focus. Deep learning models such as LogAnomaly Meng et al. (2019) and earlier sequence-learning approaches such as DeepLog Du et al. (2017) have achieved impressive results in detecting operational anomalies from large-scale unstructured logs. These methods typically learn normal system behavior from historical data and flag statistically or semantically unusual patterns as potential incidents, offering a scalable alternative to manual log inspection.
Other branches of AIOps research focus on predictive analytics and resource management. Time-series forecasting models, such as recurrent neural networks applied to cloud workload prediction Zi (2024), have been used to support proactive scaling and performance optimization. Unsupervised and explainable fault diagnosis methods further enhance interpretability in dynamic systems Chen et al. (2022). Despite these advances, existing machine learning approaches typically operate in isolation—handling logs, metrics, or traces separately. They lack a holistic, context-aware understanding of system states, and most models fail to integrate reasoning, memory, and action planning capabilities necessary for truly autonomous IT operations Abid et al. (2021); He et al. (2021). As a result, they often generate high false-positive rates or require significant human intervention to validate and act on alerts, limiting their effectiveness in fast-paced, large-scale environments Hsieh et al. (2024).
2.3 Multi-Agent Systems for Automated IT Operations
Multi-agent systems (MAS) provide another promising paradigm for distributed automation. Foundational MAS literature highlights their advantages in decentralization, autonomy, and collaborative problem-solving Wooldridge (2009). In operational environments, MAS architectures have been applied to distributed monitoring and anomaly detection, allowing agents to independently observe and report on system states across heterogeneous infrastructures Khosravifar (2018). Other works explore collaborative fault diagnosis using agent coordination strategies, enabling distributed decision-making and improved fault localization Shao et al. (2021); Xu et al. (2023).
More recent research extends MAS to cloud-native and manufacturing environments. D’Aniello et al. demonstrate distributed operational management using cooperative agents in cloud-manufacturing scenarios D’Aniello et al. (2021), while adaptive resource management studies show how MAS can respond to dynamic workloads Seitz et al. (2021). Xu et al. further illustrate MAS applications in autonomous supply chains, emphasizing task decomposition and coordination under uncertainty Xu et al. (2024). However, existing MAS designs often suffer from limited global context awareness, inefficient knowledge sharing, and difficulties maintaining long-term memory—constraints that hinder their ability to manage large-scale, highly dynamic IT infrastructures Pulikottil et al. (2021); Li et al. (2024b).
More broadly, complex real-world systems across different domains are often modeled as large-scale graphs, where robustness, efficiency, and structural properties play a central role in system behavior. Recent studies on multilayer transportation networks, geophysical systems, and graph-based learning further highlight the importance of structure-aware modeling and robustness analysis in complex environments, reinforcing the relevance of graph-structured abstractions in intelligent operational decision-making Li et al. (2022); Jiang et al. (2024); Pan et al. (2024).
2.4 Context Management and Compression in Large Language Models
Recent advances in large language models have introduced new opportunities for intelligent operations automation, but also new challenges related to long-context management Li et al. (2024a). Long-document Transformer architectures such as Longformer Beltagy et al. (2020), Reformer Kitaev et al. (2020), and BigBird Zaheer et al. (2020) provide scalable attention mechanisms to extend context windows Song et al. (2025). However, they still face limitations when applied to continuously evolving, high-volume operational data streams.
To address context overload, several context compression and sparsification techniques have been proposed. LLMLingua Jiang et al. (2023) compresses prompts to accelerate inference while preserving essential semantics, whereas contextual sparsity approaches such as DejaVu prune redundant information during inference Liu et al. (2023). Complementary memory-augmented models, such as the Recurrent Memory Transformer (RMT) Bulatov et al. (2022), introduce persistent external memory to support long-range reasoning. Recent studies further reveal that LLMs exhibit position-dependent performance degradation in long contexts ("lost in the middle") Liu et al. (2024); Ni et al. (2024), and struggle with in-context exploration without explicit guidance Krishnamurthy et al. (2024); Peng et al. (2024); Ni et al. (2025c).
While these methods improve efficiency and scalability, they are not designed specifically for IT operations environments, where context importance varies significantly across logs, metrics, events, and incidents Beyer et al. (2016). Existing LLM context mechanisms lack adaptive prioritization and domain-specific reasoning capabilities needed for complex operational decision-making Lewis et al. (2020); Borji (2023); Liu et al. (2025). Emerging hierarchical memory architectures like HiAgent Hu et al. (2025) and retrieval-augmented in-context learning Chen et al. (2024); Yu et al. (2025a) offer promising directions, yet remain unexplored in AIOps settings Syed and Anazagasty (2024); Diaz-De-Arcaya et al. (2023).
3 Theoretical Foundations
We establish the theoretical foundations of the AOI framework through formal analysis of multi-agent coordination, dynamic scheduling convergence, and context compression stability.
3.1 Markov Decision Process Formalization
We model the AOI multi-agent coordination as a Decentralized Partially Observable Markov Decision Process (Dec-POMDP), defined by the tuple $\mathcal{M}=\langle\mathcal{N},\mathcal{S},\{\mathcal{A}_{i}\},\mathcal{T},\mathcal{R},\{\Omega_{i}\},\mathcal{O},\gamma\rangle$, where:
  • $\mathcal{N}=\{Observer,Probe,Executor\}$ is the set of agents
  • $\mathcal{S}$ is the joint state space of the IT infrastructure
  • $\mathcal{A}_{i}$ is the action space for agent $i$
  • $\mathcal{T}:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\rightarrow[0,1]$ is the state transition function
  • $\mathcal{R}:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$ is the shared reward function
  • $\Omega_{i}$ is the observation space for agent $i$
  • $\mathcal{O}:\mathcal{S}\times\mathcal{A}\times\Omega\rightarrow[0,1]$ is the observation function
  • $\gamma\in[0,1)$ is the discount factor
The optimal joint policy $\pi^{*}=\{\pi_{i}^{*}\}_{i\in\mathcal{N}}$ maximizes the expected cumulative discounted reward:
$\pi^{*}=\arg\max_{\pi}\mathbb{E}\left[\sum_{t=0}^{\infty}\gamma^{t}\mathcal{R}(s_{t},a_{t})\mid\pi\right]$ (1)
3.2 Dynamic Scheduling Convergence Analysis
We prove the convergence of our dynamic scheduling strategy using potential function analysis. Define the scheduling state at time $t$ as $\sigma_{t}=(q_{t},c_{t},h_{t})$, where $q_{t}$ is the task queue state, $c_{t}$ is the compressed context, and $h_{t}$ is the historical pattern embedding.
Theorem 1 (Scheduling Convergence).
Under the AOI dynamic scheduling strategy with exploration-exploitation parameter $\lambda\in(0,1)$, the expected task completion time converges to within a factor of $(1+\epsilon)$ of the optimal static scheduling policy for any $\epsilon>0$, with probability at least $1-\delta$, provided the number of scheduling iterations $T\geq\frac{|\mathcal{A}|\log(|\mathcal{A}|/\delta)}{\epsilon^{2}\lambda^{2}}$.
Proof.
Define the potential function $\Phi(\sigma_{t})=\sum_{\tau\in q_{t}}w(\tau)\cdot d(\tau,t)$, where $w(\tau)$ is the priority weight and $d(\tau,t)$ is the delay cost. Under our scheduling policy:
$\mathbb{E}[\Phi(\sigma_{t+1})-\Phi(\sigma_{t})\mid\sigma_{t}]\leq-\eta\cdot\|\nabla_{\sigma}\Phi(\sigma_{t})\|^{2}+\frac{\lambda^{2}L^{2}}{2}$ (2)
where $\eta>0$ is the learning rate and $L$ is the Lipschitz constant of the reward function. By applying the Robbins-Monro stochastic approximation theorem, the scheduling policy converges to a stationary point. Combined with the UCB-style exploration bonus, regret analysis yields the stated bound. ∎
3.3 Context Compression Stability
We analyze the stability of the LLM-based context compression through Lyapunov analysis. Let $\mathcal{C}(t)$ denote the context state at time $t$, with compression operator $\mathcal{F}:\mathcal{C}\rightarrow\mathcal{C}_{compressed}$.
Theorem 2 (Compression Stability).
The context compression mechanism with sliding window size $w$ and overlap ratio $\rho\in(0.5,1)$ satisfies Lyapunov stability, i.e., there exists a Lyapunov function $V:\mathcal{C}\rightarrow\mathbb{R}_{\geq 0}$ such that:
$V(\mathcal{F}(\mathcal{C}))-V(\mathcal{C})\leq-\kappa\cdot\|\mathcal{C}-\mathcal{C}^{*}\|^{2}$ (3)
for some $\kappa>0$ and equilibrium state $\mathcal{C}^{*}$, ensuring bounded information loss.
Proof.
Define $V(\mathcal{C})=\frac{1}{2}\|\mathcal{C}-\mathcal{C}^{*}\|_{\mathcal{H}}^{2}$, where $\|\cdot\|_{\mathcal{H}}$ is the norm in the reproducing kernel Hilbert space induced by the LLM embedding. The compression operator satisfies:
$\|\mathcal{F}(\mathcal{C})-\mathcal{C}^{*}\|_{\mathcal{H}}\leq(1-\rho)\|\mathcal{C}-\mathcal{C}^{*}\|_{\mathcal{H}}+\epsilon_{LLM}$ (4)
where $\epsilon_{LLM}$ is the LLM approximation error bounded by model capacity. For $\rho>0.5$, contraction holds and stability follows. ∎
3.4 Information Preservation Bound
We establish theoretical guarantees on information preservation during compression.
Theorem 3 (Information Preservation).
Let $I(\mathcal{C};\mathcal{Y})$ denote the mutual information between context $\mathcal{C}$ and operational outcome $\mathcal{Y}$. The compressed context $\mathcal{C}_{comp}=\mathcal{F}(\mathcal{C})$ satisfies:
$I(\mathcal{C}_{comp};\mathcal{Y})\geq(1-\epsilon_{info})\cdot I(\mathcal{C};\mathcal{Y})$ (5)
where $\epsilon_{info}=O\left(\frac{1}{w\cdot\rho}\right)$ depends on window size $w$ and overlap ratio $\rho$.
This bound theoretically justifies our empirical observation that larger window sizes improve information preservation (Table 4).
4 Methodology
4.1 Problem statement and framework overview
Previous researches on automated IT operations have predominantly focused on rule-based execution, isolated anomaly detection, or single-agent decision-making, which fail to address the core challenges posed by modern cloud-native infrastructures—exponential growth of operational data, distributed task coordination complexity, and loss of contextual continuity in dynamic environments. These limitations lead to inefficient information processing, delayed fault remediation, and inadequate adaptability to evolving system states, hindering the realization of truly autonomous IT operations. Therefore, this paper aims to develop a scalable, context-aware multi-agent collaborative framework (AOI) that enables intelligent IT operations with minimal human intervention. However, we encounter four major challenges in practice:
  • For information overload: how to efficiently compress massive operational context while preserving critical information related to fault diagnosis and remediation.
  • For task coordination: how to dynamically schedule distributed probing and execution tasks to balance information gain and remediation efficiency based on real-time system states.
  • For context retention: how to design a memory architecture that supports efficient storage, retrieval, and reuse of operational context across multi-incident and long-running operations.
  • For operational safety: how to ensure that automated system modifications avoid destructive actions and enable rapid rollback in case of execution failures.
To tackle these challenges, we designed an innovative multi-agent framework integrated with large language model (LLM) reasoning and hierarchical memory management. Our solution consists of the following components: Multi-agent collaborative architecture- This module comprises three specialized agents—Observer for task decomposition and coordination, Probe for safe read-only information gathering, and Executor for risk-controlled system modification—ensuring clear responsibility separation and operational safety. LLM-based context compressor- This module employs a sliding-window mechanism and domain-aware summarization to intelligently compress operational context, achieving high compression ratios while retaining critical diagnostic and remediation information. Dynamic task scheduling strategy- This module adaptively prioritizes probing and execution tasks based on real-time system states, balancing exploration (information collection) and exploitation (fault remediation) to optimize mean time to resolution (MTTR) and task success rate. Three-layer memory architecture- This module comprises Raw Context Storage for unprocessed data, Task Queue Management for structured execution, and a Compressed Context Cache for LLM-processed strategic information, enabling efficient multi-scale context management.
4.2 Problem Formulation
[…]
4.3 AOI Framework Architecture
Figure 1 shows the AOI architecture, comprising 4 core components and a 3-layer memory system.
4.3.1 Observer Agent
The Observer Agent serves as the central coordination unit responsible for task analysis, decomposition, and strategic decision-making. Its core functionalities include: Task Decomposition Algorithm: Given a complex task $\tau$, the Observer employs a hierarchical decomposition strategy shown in Algorithm 1.
Algorithm 1 Dynamic Task Decomposition
0: Task $\tau$, System State $\mathbf{X}(t)$, Context $C_{compressed}$
0: Subtask Queue $\mathcal{Q}_{subtasks}$
1: Initialize $\mathcal{Q}_{subtasks}\leftarrow\emptyset$
2: Analyze task complexity $\mathcal{C}(\tau)$
3:if $\mathcal{C}(\tau)>\theta_{complex}$ then
4:  Decompose $\tau$ into atomic subtasks $\{\tau_{1},\tau_{2},...,\tau_{m}\}$
5:  for each $\tau_{i}$ in decomposed tasks do
6:   Determine task type: $type(\tau_{i})\in\{probe,execute\}$
7:   Estimate resource requirements $R(\tau_{i})$
8:   Assign priority $P(\tau_{i})$ based on dependency analysis
9:   Enqueue $(\tau_{i},type(\tau_{i}),R(\tau_{i}),P(\tau_{i}))$ to $\mathcal{Q}_{subtasks}$
10:  end for
11:else
12:  Enqueue $\tau$ directly to $\mathcal{Q}_{subtasks}$
13:end if
14:return $\mathcal{Q}_{subtasks}$
Dynamic Scheduling Strategy: The Observer implements an adaptive scheduling algorithm that balances exploration (probing) and exploitation (execution) based on current information sufficiency:
$S(t)=\arg\max_{a\in\{\mathrm{probe},\mathrm{execute}\}}\mathbb{E}[\mathrm{Reward}(a,\mathbf{X}(t),C(t))]$
where S means Schedule, the expected reward function considers information gain for probing actions and progress toward goal state for execution actions. Specifically, the reward for a probe action is quantified by the expected reduction in system state uncertainty (information gain), while the reward for an execute action is based on the estimated progress towards the target system state $\mathbf{X}^{*}$.
4.3.2 Probe Agent
The Probe Agent is a strictly read-only component for safe system information gathering. Safety-by-Design: It follows the least-privilege principle to prevent unintended modifications during diagnostics. Safety Constraints: The agent is limited to non-destructive query and read operations (e.g., database introspection and log/file reads), while any state-changing actions (e.g., writes, deletions, schema changes, or destructive commands) are explicitly prohibited. File access remains read-only, and network interactions are restricted to query-type requests.
Information Collection Strategy: The Probe employs a structured probing strategy (Algorithm 2).
Algorithm 2 Safe Information Probing
0: Probe Task $\tau_{probe}$, Target System $\mathcal{S}$
0: Raw Information $I_{raw}$
1: Initialize $I_{raw}\leftarrow\emptyset$
2: Generate probe script $script\leftarrow GenerateScript(\tau_{probe})$
3: Validate script safety: $ValidateSafety(script)$
4:if script is safe then
5:  for each command cmdcmd in script do
6:   Execute $result\leftarrow ExecuteReadOnly(cmd,\mathcal{S})$
7:   Append resultresult to $I_{raw}$
8:   if error occurred then
9:    Log error and continue with next command
10:   end if
11:  end for
12:else
13:  Reject unsafe script and request revision
14:end if
15: Store $I_{raw}$ in Memory-Raw Context
16:return $I_{raw}$
4.3.3 Executor Agent
The Executor Agent implements system modifications to achieve target operational outcomes. Unlike the Probe Agent, it can modify system state under coordination with the Observer. Conservative Execution Strategy: The Executor adopts a cautiously optimistic approach. Each execution is preceded by a full-system checkpoint, enabling instant rollback upon unexpected or critical failures, thereby supporting a fail-fast, recover-quickly paradigm (Algorithm 3).
Algorithm 3 Safe System Execution
0: Execution Task $\tau_{exec}$, Target System $\mathcal{S}$
0: Execution Result $R_{exec}$
1: Create checkpoint $checkpoint\leftarrow CreateCheckpoint(\mathcal{S})$
2: Generate execution plan $plan\leftarrow GeneratePlan(\tau_{exec})$ {Plan generation includes pre-execution validation and dependency checks}
3:for each action $a$ in plan do
4:  if RequiresSystemInfo(a) then
5:   Query Probe Agent for current state
6:   Update action parameters based on current state
7:  end if
8:  Execute $result\leftarrow ExecuteAction(a,\mathcal{S})$
9:  Validate execution result
10:  if critical failure detected then
11:   Initiate rollback to checkpointcheckpoint
12:   BREAK
13:  end if
14:end for
15: Store execution results in Memory-Raw Context
16:return $R_{exec}$
4.3.4 Context Compressor
The Context Compressor leverages large language models to intelligently process and compress operational context while preserving critical information. Semantic-Aware Compression: Unlike traditional keyword-based or statistical compression methods, our LLM-based compressor leverages the model’s semantic understanding capabilities to identify and preserve operationally critical information (e.g., error patterns, resource thresholds, causal relationships) while discarding redundant or low-signal data. Sliding Window Compression: We implement a sliding window mechanism with 50% overlap to ensure context continuity and prevent the loss of critical information that might span the boundary between two consecutive windows: $\mathrm{Window_{i}}=[start_{i},start_{i}+window_{size}]$ where $start_{i}=i\times(window_{size}\times overlap_{ratio})$ and $overlap_{ratio}=0.5$. Intelligent Compression Strategy: The compression algorithm prioritizes information based on criticality:
Algorithm 4 LLM-based Context Compression
0: Raw Context $C_{raw}$, Compression Target Ratio $r_{target}$
0: Compressed Context $C_{compressed}$
1: Split $C_{raw}$ into sliding windows $\{W_{1},W_{2},...,W_{k}\}$
2: Initialize $C_{compressed}\leftarrow\emptyset$
3:for each window $W_{i}$ do
4:  Identify critical information: $I_{critical}\leftarrow\mathrm{ExtractCritical}(W_{i})$ {Using LLM to extract fault signatures, error codes, and performance anomalies}
5:  Generate summary: $S_{i}\leftarrow\mathrm{LLMSummarize}(W_{i},I_{critical})$
6:  Append $S_{i}$ to $C_{compressed}$
7:end for
8: Merge overlapping summaries: $C_{merged}\leftarrow\mathrm{MergeOverlaps}(C_{compressed})$
9: Apply secondary compression if needed to meet $r_{target}$
10:return $C_{merged}$
4.4 Three-Layer Memory Architecture
Our memory management system consists of three specialized layers designed following the principle of data lifecycle management, where information flows from high-volume, short-term raw storage to condensed, long-term strategic memory, optimizing both retrieval speed and storage efficiency.
Layer 1: Raw Context Storage — Stores unprocessed results from Probe and Executor agents with 24-hour retention. (Feeds into Context Compressor). Layer 2: Task Queue Management — Maintains structured execution instructions with lifecycle management. (Managed by Observer and executed by Executor). Layer 3: Compressed Context Cache — Houses LLM-processed information with 7-day retention for strategic decision-making. (Populated by Context Compressor and queried by Observer).
4.5 Complexity Analysis
The computational complexity of the AOI framework is primarily determined by the LLM-based context compression module, which exhibits a time complexity of $O(n\cdot w)$, where $n$ denotes the context length and $w$ represents the compression window size. The space complexity is $O(k\cdot w)$, with $k$ corresponding to the number of overlapping sliding windows. Benefiting from the distributed multi-agent design, the framework achieves near-linear scalability with respect to the number of concurrent tasks, as each agent operates independently with minimal inter-agent communication overhead.
5 Experiments
5.1 Experimental Setup
All experiments were conducted in a controlled cloud-based environment designed to emulate realistic large-scale IT operations, with the infrastructure carefully configured to ensure reproducibility and enable controlled fault injection for stress testing. The experimental cluster consisted of eight AWS EC2 c5.4xlarge instances, each equipped with 16 vCPUs and 32 GB of RAM, supported by high-IOPS persistent storage. All nodes were interconnected through a 10 Gbps network with adjustable latency injection to simulate diverse communication conditions. The software stack was built on Ubuntu 20.04 LTS, running Docker and Kubernetes for container orchestration. System monitoring and telemetry were implemented via an integrated observability framework composed of Prometheus, Grafana, and the ELK stack Prometheus Authors (2024); Grafana Labs (2024); Elastic (2024). The large language model (LLM) backend employed a fine-tuned GPT-4 API optimized for operations-related context understanding and summarization, enabling accurate interpretation and efficient processing of system logs and alerts OpenAI (2023).
5.2 Datasets and Scenarios
We evaluate AOI on two complementary datasets spanning synthetic and real-world environments. The AIOpsLab Simulation Environment contains 1,000 fault-injection scenarios covering 50 fault types, ranging from isolated failures to cascading faults, each annotated with ground-truth resolutions and post-recovery states Chen et al. (2025). To assess practical applicability, we additionally use the Loghub Benchmark, which aggregates real-world production logs from distributed storage systems, microservices, and enterprise infrastructures Zhu et al. (2023). Logs are preprocessed via anonymization, parsing, and scenario segmentation to align with AOI’s task structure. Across both datasets, evaluation focuses on four representative operational event categories: service failure recovery, performance degradation, configuration drift, and security incidents.
5.3 Baseline Methods
To comprehensively evaluate the effectiveness of the proposed AOI framework, we compared it against four representative and reproducible baseline approaches that capture the major paradigms in automated IT operations. The Rule-based Expert System (RES) Xu et al. (2020); Sarazin et al. (2021) represents traditional automation techniques relying on deterministic rules implemented through Ansible playbooks and custom operational scripts. The Traditional AIOps Pipeline (TAP) Dang et al. (2019) follows a classical machine learning architecture that integrates anomaly detection modules with predefined scripted responses for fault remediation. The Single-Agent LLM (SA-LLM) baseline Jiang et al. (2023) applies a large language model directly to operational decision-making tasks without incorporating multi-agent collaboration or contextual memory mechanisms. Finally, the Baseline Multi-Agent System (B-MAS) D’Aniello et al. (2021) extends conventional automation through agent-based coordination but omits the intelligent context compression and adaptive scheduling strategies introduced in our AOI framework.
5.4 Evaluation Metrics
We evaluate system performance through a comprehensive set of metrics: Primary metrics include the Task Success Rate (TSR), which measures the percentage of successfully completed tasks; the Mean Time to Resolution (MTTR), referring to the average duration from task initiation to completion; the Context Compression Ratio (CCR), representing the ratio of compressed to original context size; and the Information Preservation Score (IPS), assessing the retention of critical information after compression. Secondary metrics comprise the False Positive Rate (FPR), indicating the percentage of incorrect actions taken; Resource Utilization Efficiency (RUE), which reflects computational resources consumed per task; the Scalability Index (SI), measuring performance degradation as the number of concurrent tasks increases; and the System Safety Score (SSS), evaluating the system’s capability to avoid potentially harmful actions.
5.5 Implementation Details
The system is designed for reproducibility and robustness. Each agent is implemented as a containerized microservice in Python 3.9 using FastAPI for inter-agent communication FastAPI Contributors (2024). Context compression integrates the OpenAI GPT-4 API with custom prompts optimized for operational contexts, incurring an average latency of 2.3 seconds per compression OpenAI (2023). Memory management relies on a Redis Cluster for distributed storage, failover, and persistence. Operational safety is ensured via formal TLA+ specifications and runtime monitoring with custom safety rules Lamport, Leslie (2023).
6 Results
6.1 Main Experimental Results
As shown in Table 2, AOI outperforms all baselines across metrics. It achieves the highest task success rate (94.2%), improving by 34.9% and 9.0% over rule-based and best multi-agent baselines, respectively. AOI reduces MTTR to 22.1 minutes, a 34.4% reduction from the best baseline (33.7 minutes). It also attains a 72.4% context compression ratio with 92.8% information preservation, validating the LLM-based compression. Finally, AOI achieves the lowest false positive rate (3.1%) and the highest safety score (96.7%).
Table 2: Main Experimental Results Comparison
Method TSR (%) MTTR (min) CCR (%) IPS (%) FPR (%) RUE SI SSS (%)
Rule-based Expert System (RES) 67.8 54.3 9.2 0.71 0.62 78.5
Traditional AIOps Pipeline (TAP) 75.1 42.6 35.7 81.4 7.8 0.76 0.74 82.3
Single-Agent LLM (SA-LLM) 81.5 38.2 52.9 86.7 6.9 0.79 0.77 88.6
Baseline Multi-Agent System (B-MAS) 86.4 33.7 61.8 89.3 5.6 0.82 0.81 90.1
Our AOI Framework 94.2 22.1 72.4 92.8 3.1 0.89 0.93 96.7
We evaluate cross-dataset generalization on four heterogeneous benchmarks—HDFS, BGL, OpenStack, and AIOpsLab—covering both real-world and simulated operational environments. As shown in Figure 3, AOI consistently outperforms all baselines across datasets, indicating robust transferability beyond the training distribution. Quantitative results in Table 7 further confirm this trend: AOI achieves the highest Task Success Rate (TSR) on all datasets (92.1–94.2%) while simultaneously reducing Mean Time To Resolution (MTTR) to around 22 minutes, substantially outperforming the strongest baseline (B-MAS). Notably, the performance gains are stable across diverse log characteristics and system settings, demonstrating that AOI generalizes well across operational domains rather than overfitting to a specific dataset.
Figure 3: Cross-dataset performance comparison across HDFS, BGL, OpenStack, and AIOpsLab benchmarks. AOI consistently outperforms all baseline methods across diverse operational environments, demonstrating strong generalization capability.
6.2 Ablation Studies
Our ablation studies (Table 3) systematically validate the necessity of each component in the AOI framework. The absence of the context compressor results in a 5.7% TSR drop and a 7.7-minute MTTR increase, directly attributing its value to filtering noise. The absence of dynamic scheduling causes a 20.8% MTTR increase, linking its function to real-time prioritization. Replacing the three-layer memory incurs a 4.8% performance loss, confirming its role in context reuse. Ultimately, the single-agent baseline’s nearly 10% lower TSR conclusively establishes the effectiveness of our multi-agent coordination and specialization.
Table 3: Ablation Study Results
Configuration TSR (%) MTTR (min) CCR (%)
AOI (Full) 94.2 22.1 72.4
w/o Context Compressor 88.5 29.8 N/A
w/o Dynamic Scheduling 90.1 26.7 70.5
w/o Three-Layer Memory 89.4 27.5 65.2
Single-Agent Version 84.6 31.4 60.3
6.3 Parameter Sensitivity Analysis
We analyze the sensitivity of AOI to three key parameters: context compression window size, scheduling balance factor, and memory retention policy.
Compression Window Size. Table 4 and Figure 4 show that increasing the compression window size consistently improves TSR, CCR, and IPS up to a moderate range, after which gains become marginal. In particular, a window size of 768 tokens achieves the best overall trade-off, reaching a CCR of 72.4% and an IPS of 92.8%, while maintaining strong TSR. Larger windows (e.g., 1024 and 1536 tokens) provide only negligible improvements at increased computational cost, indicating diminishing returns beyond this point.
Scheduling Strategy Parameters. We vary the exploration–exploitation balance factor $\lambda\in[0.2,0.8]$ and observe that performance peaks at $\lambda=0.35$, suggesting that moderate exploration is critical for effective decision scheduling.
Memory Retention Policies. Retaining episodic memory for 72 hours improves TSR by 3.5% compared to 24-hour retention, while extending retention to 120 hours yields minimal additional gains with higher resource overhead.
Table 4: Parameter sensitivity for context compression window size (results averaged over 5 runs). IPS = Information Preservation Score.
Window size (tokens) TSR (%) CCR (%) IPS (%)
256 $89.1\pm 1.3$ $58.7$ $86.2$
512 $91.4\pm 1.0$ $66.0$ $90.1$
768 $\mathbf{94.2\pm 0.8}$ $\mathbf{72.4}$ $\mathbf{92.8}$
1024 $94.5\pm 0.9$ $74.0$ $92.9$
1536 $94.6\pm 1.0$ $74.6$ $92.7$
Figure 4: Impact of Key Parameters on Performance
6.4 Scalability Study
We evaluate the scalability of AOI under increasing levels of concurrency using the AIOpsLab simulation. As shown in Figure 5, AOI maintains stable performance as the number of concurrent tasks grows, with task success rates remaining above 88% up to 50 parallel tasks. Table 5 provides quantitative results, showing a gradual and controlled increase in MTTR as concurrency rises, while resource utilization efficiency (RUE) scales near-linearly with task load. These results indicate that AOI avoids performance collapse under high concurrency and can support enterprise-scale deployments with sustained effectiveness.
Table 5: Scalability under increasing concurrent tasks (AIOpsLab simulation). Values are mean over 5 runs.
Tasks TSR MTTR RUE
(%) (min) (CPU-s/task)
1 $95.3\pm 0.6$ $21.6\pm 0.9$ $123$
5 $94.6\pm 0.7$ $22.3\pm 1.4$ $126$
10 $93.8\pm 0.8$ $23.5\pm 1.3$ $131$
20 $92.1\pm 1.2$ $25.8\pm 1.4$ $157$
50 $88.9\pm 1.6$ $31.2\pm 2.7$ $194$
Figure 5: Scalability analysis of the AOI framework under increasing concurrent task loads. The framework maintains high task success rates ($88\%$) and near-linear resource utilization efficiency up to 50 concurrent tasks, demonstrating robust scalability for enterprise-scale deployments.
6.5 Qualitative Analysis and Case Studies
Case Study 1: Complex Cascading Failure. In a simulated e-commerce infrastructure, AOI identified a cascading database connection failure within 4.3 minutes, automatically triggered dependency recovery, and fully restored service in 18.5 minutes. Competing systems required over 35 minutes and often failed to detect intermediate dependencies.
Case Study 2: Configuration Drift Resolution. In a Kubernetes-based microservice deployment, AOI detected a policy mismatch between ingress and service configurations. The Probe agent safely confirmed the issue, and the Executor executed a validated rollback plan with zero packet loss, resulting in overall downtime of less than 3 minutes.
Failure Case Analysis. AOI exhibited limitations in scenarios involving extremely rapid failure propagation (less than 2 seconds), unseen hybrid fault types absent from the training corpus, and situations where GPT-based compression latency caused temporary decision delays under extreme load. These cases highlight areas for improvement and guide ongoing efforts toward online learning and real-time adaptive compression.
Table 6: Ablation study (5-run mean $\pm$ std). Each row disables specified component(s).
Configuration TSR MTTR CCR
(%) (min) (%)
AOI (Full) $\mathbf{94.2\pm 0.8}$ $\mathbf{22.1\pm 1.0}$ $\mathbf{72.4}$
w/o Context Compressor $88.5\pm 1.2$ $29.8\pm 1.5$ N/A
w/o Dynamic Scheduling $90.1\pm 1.0$ $26.7\pm 1.3$ $70.5$
w/o Three-Layer Memory $89.4\pm 1.1$ $27.5\pm 1.4$ $65.2$
w/o Safety Verifier $92.0\pm 0.9$ $24.3\pm 1.2$ $72.1$
w/o Checkpointing $91.1\pm 1.0$ $25.8\pm 1.6$ $71.9$
Single-Agent Version $84.6\pm 1.4$ $31.4\pm 1.9$ $60.3$
Table 7: Cross-dataset summary: Task Success Rate (TSR) and Mean Time To Resolution (MTTR). Results are reported as mean $\pm$ std over 5 independent runs.
Dataset Method TSR (%) MTTR (min) Dataset AOI TSR / MTTR
HDFS RES $62.4\pm 2.1$ $58.7\pm 3.5$ HDFS (AOI) $\mathbf{92.1\pm 1.1}$ / $\mathbf{22.9\pm 1.2}$
TAP $71.8\pm 1.9$ $44.9\pm 2.8$ B-MAS (best) $84.7\pm 1.4$ / $34.8\pm 1.9$
SA-LLM $78.9\pm 1.6$ $40.6\pm 2.1$
B-MAS $84.7\pm 1.4$ $34.8\pm 1.9$
AOI $\mathbf{92.1\pm 1.1}$ $\mathbf{22.9\pm 1.2}$
BGL RES $66.1\pm 2.4$ $55.0\pm 3.8$ BGL (AOI) $\mathbf{93.4\pm 0.9}$ / $\mathbf{21.7\pm 1.0}$
TAP $74.0\pm 2.0$ $43.7\pm 3.0$ B-MAS (best) $85.6\pm 1.3$ / $33.2\pm 2.0$
SA-LLM $80.3\pm 1.7$ $39.1\pm 2.4$
B-MAS $85.6\pm 1.3$ $33.2\pm 2.0$
AOI $\mathbf{93.4\pm 0.9}$ $\mathbf{21.7\pm 1.0}$
OpenStack RES $64.0\pm 2.3$ $56.9\pm 3.6$ OpenStack (AOI) $\mathbf{92.7\pm 1.0}$ / $\mathbf{22.4\pm 1.3}$
TAP $73.6\pm 2.1$ $45.5\pm 2.9$ B-MAS (best) $86.1\pm 1.2$ / $34.1\pm 1.8$
SA-LLM $79.8\pm 1.8$ $39.8\pm 2.5$
B-MAS $86.1\pm 1.2$ $34.1\pm 1.8$
AOI $\mathbf{92.7\pm 1.0}$ $\mathbf{22.4\pm 1.3}$
AIOpsLab (sim) RES $67.8\pm 2.0$ $54.3\pm 3.3$ AIOpsLab (AOI) $\mathbf{94.2\pm 0.8}$ / $\mathbf{22.1\pm 1.0}$
TAP $75.1\pm 1.8$ $42.6\pm 2.7$ B-MAS (best) $86.4\pm 1.1$ / $33.7\pm 1.7$
SA-LLM $81.5\pm 1.5$ $38.2\pm 2.2$
B-MAS $86.4\pm 1.1$ $33.7\pm 1.7$
AOI $\mathbf{94.2\pm 0.8}$ $\mathbf{22.1\pm 1.0}$
7 Discussion
7.1 Key Insights and Implications
Our experimental evaluation reveals several important insights about multi-agent collaborative frameworks for IT operations:
Context Management is Critical: The significant performance difference between configurations with and without intelligent context compression demonstrates that information overload is a major bottleneck in automated operations. Our LLM-based compression approach successfully addresses this challenge while preserving critical operational details.
Agent Specialization Improves Safety: The clear separation of responsibilities between read-only Probe agents and write-capable Executor agents provides inherent safety guarantees that monolithic approaches cannot achieve. This design principle should be adopted more widely in automation frameworks.
Dynamic Scheduling Outperforms Static Approaches: Our adaptive scheduling strategy consistently outperforms fixed exploration-exploitation ratios, suggesting that operational contexts require sophisticated decision-making that adapts to current system state and historical patterns.
7.2 Advantages and Limitations
The AOI framework offers several notable advantages. Its distributed agent architecture ensures scalability, allowing performance to grow linearly with system complexity. Built-in safety mechanisms enhance safety by preventing destructive operations, while dynamic scheduling provides adaptability, effectively responding to varying operational conditions. Additionally, intelligent context compression improves efficiency by reducing information overload. Despite these strengths, the framework has certain limitations. Its reliance on external LLM services introduces latency and potential service dependencies. Performance is also contingent on the quality and coverage of training scenarios, and the processes of context compression and multi-agent coordination incur additional computational overhead. Finally, the system exhibits limited capability in handling completely unprecedented failure scenarios, representing novel failure modes that remain challenging.
7.3 Practical Deployment Considerations
Organizations considering adoption of the AOI framework should account for several practical factors. First, infrastructure requirements include container orchestration capabilities and sufficient computational resources to support LLM operations, which may increase infrastructure costs. Second, integration challenges arise as existing monitoring and automation tools must interface with the agent framework, making API compatibility and data format standardization essential. Third, staff training is necessary, as operations teams need to understand multi-agent system concepts and framework-specific procedures, requiring careful change management. Finally, in regulated industries, regulatory compliance must be ensured, with automated decision-making aligned to relevant requirements and accompanied by appropriate audit trails.
8 Conclusion
This paper presents AOI, a multi-agent framework for intelligent IT operations that addresses key challenges in modern infrastructure management. AOI integrates a specialized three-agent architecture, LLM-based context compression, adaptive task scheduling, and hierarchical memory to enable safe, efficient, and scalable operations. Extensive experiments show that AOI significantly improves task success, resolution time, and overall operational efficiency over existing approaches, while ablation and case studies further validate the effectiveness of each component in realistic scenarios.
Limitations
Despite the mentioned strengths, the framework has certain limitations. Its reliance on external LLM services introduces latency and potential service dependencies. Performance is also contingent on the quality and coverage of training scenarios, and the processes of context compression and multi-agent coordination incur additional computational overhead. Additionally, the system exhibits limited capability in handling completely unprecedented failure scenarios, representing novel failure modes that remain challenging.
Appendix A Future Work
Our research also highlights the importance of safety-by-design principles in automated operations systems. The clear separation between information gathering and system modification agents provides inherent safety guarantees that should be considered essential in any operational automation framework.
Several promising research directions emerge from this work. Federated Learning Integration could enable AOI frameworks across multiple organizations to share operational knowledge while preserving privacy and security. Advanced Context Compression can be explored by developing domain-specific models trained on operational data to further improve compression efficiency and information retention. Formal Verification Extensions offer the potential to provide stronger guarantees regarding system behavior and safety properties. Human-AI Collaboration could be enhanced through interfaces and protocols that support seamless interaction in complex operational scenarios requiring human judgment. Finally, Real-time Adaptation would allow the framework to respond to rapidly changing operational conditions via online learning and real-time model updates. As modern IT infrastructure grows increasingly complex, intelligent automation becomes not only beneficial but essential. The AOI framework lays a foundation for next-generation operational automation systems capable of managing this complexity while ensuring safety and reliability. With the ongoing adoption of cloud-native architectures and microservices, frameworks like AOI will become critical enablers of scalable, resilient, and efficient operations.